OpenStudy (anonymous):

How would I convert .0044344mol/L (mean concentration of Fe sample) to mg/ml

OpenStudy (anonymous):

\[\frac{ 0.0044344molesFe }{ 1liter } \times \frac{ 55.85gFe }{ 1mole } \times \frac{ 1000mg }{ 1g } \times \frac{ 1liter }{ 1000ml }=0.24766124 mg/mL\]